# figurec (eccv) Package eccv Warning: Package ‘hyperref’ is loaded with option ‘pagebackref’, which is *not* recommended for camera-ready version

### v1

@@ 0.C.2 Architecture
Figure 10 describes the model architecture in more detail. It is based on SVD [5], which in turn is based on Video LDM [6], modified for camera pose conditioning. The $T=14$ CLIP embeddings are fed to the network via multiple spatial and temporal cross-attention blocks throughout the network. Separately, the micro-conditioning mechanism takes place to pass the embeddings of the diffusion timestep, frame rate, camera transformation, motion bucket value, and conditioning augmentation strength to the network by summing it together with feature channels at various residual blocks placed throughout the network, with additional linear projections in-between to accommodate varying embedding sizes. Concretely, assuming the camera always looks at the same location in 3D space for simplicity, [^1] This is $(0,0,1)$, i.e. 1m above the center of the ground plane, in Kubric-4D. the relative extrinsics matrix $\Delta\mathcal{E}$ is parameterized as $(\Delta\phi,\Delta\theta,\Delta r)$. The angles are subsequently encoded with Fourier positional encoding before being embedded through an MLP. Note that the input camera poses are not required to be known – only the desired relative transformation should be given.
Figure 10: Network architecture. Our model performs diffusion in latent space [49, 33]. The input video is encoded by a VAE, and then channel-concatenated with the noisy sample. At training time, the output video is estimated and supervised; at inference time, multiple denoising steps are performed. In both cases, per-frame CLIP embeddings and other relevant pieces of information (frame rate, desired camera pose transformation, and motion value) condition the U-Net in different ways.
@@ 0.C.3 Data and training
In Kubric-4D, pairs of input and output video clips are always temporally synchronized, but with $T=14$ frame indices sampled randomly within the 60 available frames from the dataset. The original FPS is 24, and since the frame stride is randomly uniformly sampled within $[1,4]$, the actual FPS when finetuning therefore varies between 6 and 24. In ParallelDomain-4D, each scene has 50 frames available at 10 FPS, from which we randomly subsample clips but only at a frame stride of either 1 or 2, which implies FPS values between 5 and 10.
In Kubric-4D, the camera pose respects the following bounds (both across time and across input/output) with respect to the spherical coordinate system: azimuth angle $\phi_{1\ldots T}\in[0\degree,360\degree]$, elevation angle $\theta_{1\ldots T}\in[0\degree,50\degree]$, radial distance $r_{1\ldots T}\in[12,18]$.222 Since the dataset is synthetic and the radius $r$ does not have an inherent meaning, it is worth nothing that the average diameter of an object is 1.88m, and that all objects are randomly spawned within these bounds in Euclidean coordinates: $x\in[-7,7],y\in[-7,7],z\in[0,7]$ (where Z is up). The target camera pose transformation for the default model (max 90°) has a limited maximum transformation “strength” in the sense that from start to end, the azimuth, elevation, and radius all vary within the following bounds: $|\Delta\phi|\leq 90\degree,|\Delta\theta|\leq 30\degree,|\Delta r|\leq 3$. The horizontal field of view is 53.1° everywhere.
@@
The trajectories are typically uniformly sampled, except for the elevation angle $\theta$; in this case, uniform sampling for the starting point happens in terms of $\sin\theta$ instead of the angle $\theta$ directly. This is done in order to ensure an equal spread over (i.e. a uniform distribution on the surface of) the (relevant subset of the) unit sphere. The input camera extrinsics $\mathcal{E}_{src}$ is static, and the output camera extrinsics $\mathcal{E}_{dst,t}$ interpolates linearly over time in spherical coordinate space.
In ParallelDomain-4D, the source viewpoint is a forward-facing camera mounted on the virtual ego car at a fixed position of $(1.6,0,1.55)$ in 3D world space, where X points forward and Z points up. For simplicity, the camera pose is not controllable in the experiments described in our paper – instead, the destination viewpoint is fixed at $(-8,0,8)$, looking forward and down at $(5.6,0,1.55)$. To maximize the temporal smoothness of the generated video, the camera trajectory is interpolated in Euclidean space, not linearly but rather according to a sine wave function, i.e. following $\frac{1-\cos\left(\frac{t}{\pi(T-1)}\right)}{2}$, assuming $t$ increases step-wise from 0 to $T-1$. The horizontal field of view is 85° everywhere.
@@ 0.C.5 Evaluation
For each dataset separately, all models and all variants are evaluated on the same test split. For each scene, we randomly sample a subclip within the available video with $T=14$ frames and a variable frame rate chosen within the same range as during training time. Then, for Kubric-4D, four different target camera poses (with angles up to azimuth ±$\pm 90\degree$ for Kubric-4D) are randomly. To encourage moderately difficult input videos with higher than average degrees of occlusion, we set the starting elevation angle to be always $\theta_{1}=5\degree$, but all other angles are chosen randomly within the same ranges as during training. These randomization parameters at test time are only chosen once and then fixed across all evaluation experiments. We let probabilistic (i.e. diffusion) models (Ours, Vanilla SVD, ZeroNVS) generate four samples for each of these trajectories, averaging results, but the other methods (HexPlane, 4D-GS, DynIBaR) are only executed once for each scene and for each set of output camera angles.

### v2

@@ 0.C.2 Architecture
Figure 10 describes the model architecture in more detail. It is based on SVD [5], which in turn is based on Video LDM [6], modified for camera pose conditioning. The $T=14$ CLIP embeddings are fed to the network via multiple spatial (S-Attn) and temporal (T-Attn) cross-attention blocks throughout the network. Separately, the micro-conditioning mechanism takes place to pass the embeddings of the diffusion timestep, frame rate, camera transformation, motion bucket value, and conditioning augmentation strength to the network by summing it together with feature channels at various residual blocks placed throughout the network, with additional linear projections in-between to accommodate varying embedding sizes. Concretely, assuming the camera always looks at the same location in 3D space for simplicity, [^2] This is $(0,0,1)$, i.e. 1m above the center of the ground plane, in Kubric-4D. the relative extrinsics matrix $\Delta\mathcal{E}$ is parameterized and given as $(\Delta\phi,\Delta\theta,\Delta r)$. The angles are subsequently encoded with Fourier positional encoding before being embedded through an MLP. Note that the input camera poses are not required to be known – only the desired relative transformation should be given.
Figure 10: Network architecture. Our model performs diffusion in latent space [52, 35]. The input video is encoded by a KL-VAE, and then channel-concatenated with the noisy sample. At training time, the output video is estimated and supervised; at inference time, multiple denoising steps are performed. In both cases, per-frame CLIP embeddings condition the U-Net by means of cross-attention, and and other relevant pieces of information (frame rate, desired camera pose transformation, and motion value) condition the U-Net by adding their embeddings onto the feature vectors in-between convolutions.
@@ 0.C.3 Data and training
In Kubric-4D, pairs of input and output video clips are always temporally synchronized, but with $T=14$ frame indices sampled randomly within the 60 available frames from the dataset. The original FPS is 24, and since the frame stride is randomly uniformly sampled among $\{1,2,3,4\}$, the actual FPS when finetuning therefore belongs to $\{6,8,12,24\}$. In ParallelDomain-4D, each scene has 50 frames available at 10 FPS, from which we randomly subsample clips but only at a frame stride in $\{1,2\}$ determined by a coin flip, which implies an FPS value in $\{5,10\}$.
In Kubric-4D, the camera pose $\mathcal{P}=(\phi,\theta,r)$ respects the following bounds (both across time and across input/output) with respect to the spherical coordinate system: azimuth angle $\phi_{1\ldots T}\in[0\degree,360\degree]$, elevation angle $\theta_{1\ldots T}\in[0\degree,50\degree]$, radial distance $r_{1\ldots T}\in[12,18]$.333 Since the dataset is synthetic and the radius $r$ does not have an inherent meaning, it is worth nothing that the average diameter of an object is 1.88m, and that all objects are randomly spawned within these bounds in Euclidean coordinates: $x\in[-7,7],y\in[-7,7],z\in[0,7]$ (where Z is up). The target camera pose transformation for the default model (max 90°) has a limited maximum transformation “strength” in the sense that from start to end, the azimuth, elevation, and radius all vary within the following bounds: $|\Delta\phi|\leq 90\degree,|\Delta\theta|\leq 30\degree,|\Delta r|\leq 3$. The horizontal field of view is 53.1° everywhere.
@@
The trajectories are typically uniformly sampled, except for the elevation angle $\theta$; in this case, uniform sampling for the starting point happens in terms of $\sin\theta$ instead of the angle $\theta$ directly. This is done in order to ensure an equal spread over (i.e. a uniform distribution on the surface of) the (relevant subset of the) unit sphere. The input camera extrinsics $\mathcal{E}_{src,t}$ is static, and the output camera extrinsics $\mathcal{E}_{dst,t}$ interpolates linearly over time in pose description space, i.e. in spherical coordinates with $\alpha=\frac{t}{T-1}$.
In ParallelDomain-4D, the source viewpoint is a forward-facing camera mounted on the virtual ego car at a fixed position of $(1.6,0,1.55)$ in 3D world space, where X points forward and Z points up. For simplicity, the camera pose is not controllable in the experiments described in our paper – instead, the destination viewpoint is fixed at $(-8,0,8)$, looking forward and down at $(5.6,0,1.55)$. To maximize the temporal smoothness of the generated video, the camera trajectory is interpolated in Euclidean space, not linearly but rather according to a sine wave function, i.e. following $\alpha=\frac{1-\cos\left(\frac{t}{\pi(T-1)}\right)}{2}$, assuming $t$ increases step-wise from 0 to $T-1$. The horizontal field of view is 85° everywhere.
@@ 0.C.5 Evaluation
For each dataset separately, all models and all variants are evaluated on the same test split. For each scene, we randomly sample a subclip within the available video with $T=14$ frames and a variable frame rate chosen within the same range as during training time. Then, for Kubric-4D, four different target camera poses (with angles up to azimuth ±$\pm 90\degree$ for Kubric-4D) are randomly sampled once. To encourage difficult input videos with higher than average degrees of occlusion, we set the starting elevation angle to be always $\theta_{1}=5\degree$, but all other angles are chosen randomly within the same ranges as during training. These randomization parameters at test time are only chosen once and then fixed across all evaluation experiments. We let probabilistic (i.e. diffusion) models (Ours, Vanilla SVD, ZeroNVS) generate four samples for each of these trajectories, averaging results, but the other methods (HexPlane, 4D-GS, DynIBaR) are only executed once for each scene and for each set of output camera angles.